OpenStudy (anonymous):

what is the formula for finding how long a line is? Like for example: how far away is (-3,4) from (5,4)?

OpenStudy (amistre64):

its the pythag thrm really, the distance is the hypotenuse of a right triangle

OpenStudy (amistre64):

d^2 = (x-x0)^2 + (y-y0)^2

OpenStudy (amistre64):

d^2 = (-3-5)^2 + (4-4)^2 d^2 = 64 d = sqrt(64) = 8

OpenStudy (amistre64):

the other way is to notice this is a flat line that goes from -5 to 3 and so is just 8 units long